# Multi-modal Generative AI: Multi-modal LLMs, Diffusions and the Unification

### v2

@@ II-D Video LLM
Following the success of Image LLMs, researchers start exploring the training of Video LLMs [47]. Typically, videos are viewed as sequences of image frames (some Video LLMs incorporate other modalities like audio or speech), so Video LLMs have a higher computational complexity compared to Image LLMs. The challenge of collecting high-quality video datasets further complicates the training process, making early fusion architectures computationally exhaustive. As a result, almost all the existing Video LLMs adopt the alignment architectures.
@@ II-D1 Alignment-Architecture Video LLM
The video LLM architecture is similar to that of Image LLMs with alignment architectures. By sampling a fixed number of frames or using a fixed frames-per-second (FPS) rate, videos are reduced to a limited set of images. The visual embeddings of each image are then extracted using a visual encoder. These features are sequentially concatenated in the order of the frames and connected to the LLM via an alignment module. In earlier works, VideoChat [48] utilizes a Q-former structure as the alignment module, while VideoLLaMA [49] introduces an audio encoder and an audio Q-former to handle audio signals. Video-ChatGPT [50] takes a different approach by average-pooling each frame’s patch embeddings along the spatial and temporal dimensions before using a linear layer as the alignment module. Training Video LLMs also follows an “alignment then instruction tuning” strategy. While additional GPT-annotated or human-annotated video datasets are collected, image datasets can also be leveraged by treating images as single-frame videos.
Recent successful efforts focus on improving performance by refining the alignment module and scaling up the model and dataset sizes. For instance, VideoLLaMA2 [51] improves the alignment module to model the connections across temporal and spatial dimensions. It also gathers datasets for tasks such as captioning, classification, and question answering. Qwen2.5-VL [52] and InternVL3 [53] leverage diverse training data including images, videos, and interleaved image–text pairs to build powerful vision-language models.
@@ II-D2 Challenges and Limitations in Video LLM
Despite the advancements in Video LLMs, nearly all existing models rely on sampling frames and encoding them individually through image encoders. This approach may be favored due to several reasons: image encoders are less computationally intensive compared to video encoders, they offer better alignment with textual data, and they facilitate unification with Image LLMs. However, this methodology comes with a significant limitation. Specifically, the process of sampling frames can lead to the complete loss of information that occurs between sampled frames. As a result, these models fail to capture the continuous motion and trajectories of objects, which are essential for understanding dynamic scenes and activities within a video.
Now we have discussed the multi-modal LLM for visual understanding. Next, we will discuss another important topic of multi-modal generative AI, i.e., multi-modal diffusion models for visual generation.
@@ III Multi-modal Diffusion for Generation
Figure 4: The comparison of fundamental architectures of GANs, VAEs, and diffusions.
@@ III-A Preliminaries
We will first introduce some preliminaries, including traditional generative models, i.e., generative adversarial networks (GANs) and Variational AutoEncoders (VAEs). Then the diffusion probabilistic modeling will be introduced, followed by comparisons of GANs, VAEs and Diffusions in Fig. 4.
@@ III-A1 Generative Adversarial Networks
The generative adversarial network (GAN) [61] is one of the earliest neural architectures designed to generate visual content such as images [62] and videos [63]. The main idea of GANs involves two networks: a generator $\mathcal{G}$ and a discriminator $\mathcal{D}$. Specifically, $\mathcal{G}$ aims to generate visual content from a noise vector $z$, while $\mathcal{D}$ is trained to distinguish between real visual samples $x$ and generated ones $\mathcal{G}(z)$. These two networks are trained in an adversarial manner: the generator tries to produce outputs that can fool the discriminator, and the discriminator strives to accurately classify real versus fake samples. The training process forms a min-max game, where the generator learns to generate increasingly realistic samples to deceive a progressively stronger discriminator.
@@ III-A2 Variational AutoEncoder
Variational AutoEncoder [64] (VAE) is another typical generative model. Unlike GANs, autoencoders have an encoder-decoder architecture that uses an encoder $\mathcal{E}$ to present the visual content $x$ to a latent code $z=\mathcal{E}(x)$ and a decoder $\mathcal{D}$ to reconstruct the data $\hat{x}=\mathcal{D}(z)\approx x$. However, normal autoencoders have no constraints to the latent space, which makes it overfit the dataset easily. To solve the problem, VAEs make a regularization to the latent space and sample $z$ from a distribution $p_{\theta}$, typically a Gaussian distribution, where $\theta$ is the parameters of the encoder-decoder model.
@@ III-A3 Diffusion Probabilistic Modeling
Remark. GANs, VAEs, and diffusion models are all generative models. Compared to GANs which train both the generator and discriminator, the diffusion models have explicit probabilistic modeling, and only train a denoising network $\epsilon_{\theta}$, which is more stable. Similarly, VAEs train both an encoder and a decoder. Moreover, diffusions denoise for each image $T$ times in the training phase, resulting in $T$ variants of each image as augmentation. These augmented images in turn help the denoising network to better model the data distribution $p_{\theta}(x_{0})$, leading to better generation results.
@@ III-A4 Latent Diffusion Model
Note that there is an additional input $c$ of the denoising network that is for conditional generation, e.g., as for the text-to-image generation task, $c$ could be the representation of the text prompt [68]. Also, $c$ could be other conditions, such as layout [69] or semantic maps [70].
Since most computation including the training and iterative inference is conducted in the lower-dimension latent space, the LDM model exhibits high efficiency. Therefore, most text-to-image and text-to-video models adopt the LDM structure.
Figure 5: Comparison between pixel-level diffusion models and latent diffusion models.
@@ III-B Text-to-Image Generation
III-B1 Text-to-Image Diffusion Model
As mentioned in the preliminary part, diffusion models can be broadly categorized into two branches: pixel-based and latent-based [71]. In the early development stage, the denoising process is typically applied directly in the pixel space. For instance, GLIDE [72] is a pioneering work in photorealistic image generation with text guidance, using a 3.5 billion parameter diffusion model that employs a text encoder to condition on natural language descriptions. GLIDE also explores the use of CLIP guidance and classifier-free guidance in diffusion models, finding that classifier-free guidance produces higher-quality images. Besides, Imagen [73] follows GLIDE and adopts classifier-free guidance for its pixel-based diffusion model. The key difference between them is that GLIDE trains text encoder and diffusion model together with text-image pair, while Imagen utilizes pretrained and frozen large transformer language models, leveraging their strong text understanding capabilities to enhance sample fidelity and image-text alignment.
@@ III-B1 Text-to-Image Diffusion Model
GLIDE [72], Imagen [73], Stable Diffusion [74], and DALL-E2 [75] are all pioneering works that represent different technological pathways in the field of text-to-image generation. Despite their differences, some common trends have emerged in their development. First, latent-based diffusion methods have become increasingly prevalent due to their advantages in conserving computational resources and generating high-quality images. Second, compared to classifier guidance [76], classifier-free guidance [77] are widely adopted in these works, where the label in a class-conditional diffusion model is replaced with a null label at a fixed probability during training. Third, U-Net traditionally serves as the backbone of the diffusion model, facilitating denoising and the gradual generation of high-quality images.
Despite its advantages in high-resolution image generation, U-Net’s specific structures, such as ResBlocks and convolutional operations, limit its scalability. In contrast, Transformers, which are better suited to handle larger-scale data and tasks, are emerging as strong contenders to U-Net. The Diffusion Transformer (DiT) [78] represents a class of diffusion models that replaces the commonly used U-Net backbone with a transformer backbone as shown in Fig. 6. This approach is supported by empirical findings suggesting that the U-Net inductive bias is not crucial to the performance of diffusion models. Additionally, utilizing a transformer backbone enables the diffusion model to leverage the best practices of transformers, such as architectural design and training paradigms, along with their good properties like scalability, robustness, and efficiency. Specifically, DiT adheres to the foundation of the Latent Diffusion Model (LDM) framework and emulates the design of the Vision Transformer (ViT) by introducing a comprehensive DiT design space, including patch size, transformer block architecture, and model size. The first layer of DiT, termed patchify, converts the spatial input into a sequence of tokens by linearly embedding each patch. Following the patchify step, the input tokens are processed through a sequence of transformer blocks that incorporate conditioning such as time and label. The proposed transformer design includes adaptive layer norm (adaLN) block, cross-attention block, and in-context conditioning block. After the final block, a transformer decoder translates the image tokens into output predictions. The difference between U-Net-based diffusion model and Transformer-based diffusion model is illustrated in Fig. 6.
@@
III-B2 Controllable Generation with Diffusion Model
Despite the success of diffusion models in generating photorealistic images, this text-to-image technique falls short of fully meeting the increasing and diverse user requirements, such as fine-grained control or specific customization. For instance, creating a portrait of an ordinary individual based solely on a name and physical description is beyond the capabilities of current text-to-image diffusion models. As a result, ongoing efforts are improving diffusion models to allow for more precise control through additional conditions beyond text.
Figure 8: Model-based, tuning-based, and training-free controllable text-to-image generation.
Compared to original text-based image generation, controllable generation involves introducing additional conditions $c_{add}$ without compromising the original text condition $c_{text}$. From the perspective of control mechanisms, related methods can be categorized into three classes: model-based, tuning-based, and training-free [82], as illustrated in Fig. 8. Model-based methods incorporate extra models to encode the additional conditions and integrate them into the diffusion process. For instance, InstantBooth [83] employs a patch encoder and a concept encoder to encode personalized sample images and incorporates adapter layers to the U-Net for conditions interaction. These extra models, including encoders and adapters, are trainable, while other components of the model remain frozen. Tuning-based methods do not require extra models but instead fine-tune certain parts of the original diffusion model to adapt to specific conditions. For example, Textual Inversion [84] fine-tunes the text encoder, while Dreambooth [85] fine-tunes the U-Net. In these tuning-based methods, Parameter-Efficient Fine-Tuning (PEFT) techniques are often employed to replace traditional fine-tuning, thereby reducing computational resources. Training-free methods eliminate the need for any training or fine-tuning process, instead controlling generation by leveraging the intrinsic capabilities of the U-Net structure, such as attention. For example, StyleAligned [86] achieves consistent style generation by employing minimal “attention sharing” during the diffusion process, where all images share self-attention with the reference image.
Each of these three methods has its own strengths and weaknesses. Model-based methods introduce additional models and require a tuning process, generally consuming the most computational resources. However, once the encoders and adapters are fully trained, they can be easily adapted to different conditions. Tuning-based methods save computational resources by not incorporating extra models, but they are limited to adapting to a single specific condition per fine-tuning. Training-free methods do not require any extra models or fine-tuning time, but they are restricted to controlling only a limited range of conditions, such as layout or style.
Controllable generation involves customizing various aspects of an image, such as subject, layout, and style [82]. Among these, the primary control condition is the subject of an image, known as subject-driven generation [85, 87]. For instance, when a user describes an image with a phrase like “a dog is running on the beach”, the dog may not be an arbitrary one but a specific, familiar dog. To achieve this, sample images of the dog are provided as an additional condition, and an uncommon word, such as “[V]”, “sks”, or “S*”, is assigned to the description to represent the specific subject. A related but more specialized area is person-driven generation [88, 89], which focuses on maintaining a consistent human identity, depicting a person with different expressions, postures, and actions. Compared to subject-driven generation, layout, and style-driven generation focus on the overall composition of the image. Layout conditions [90, 91] control the relative positions of different subjects and the background, while style conditions [92, 93] determine the artistic style of the image, such as oil painting, black-and-white, or line art. Additionally, other novel conditions, such as sound, brain signals, and semantic maps, are being explored to control text-to-image generation, offering new and subtle ways to influence the mood and perception of an image.
Beyond addressing specific conditions, many studies explore complex control involving multiple conditions. For example, generating multiple customized subjects performing user-defined actions is particularly challenging because the model might confuse or forget the specified conditions. Thus, multiple conditioning control involves more than simply combining various specific conditions; it requires their interaction in a well-designed manner.
Existing methods for controlling generation with multiple conditions can be categorized as follows. Joint training methods [94, 95] rely on multi-condition encoders and specialized training strategies to manage diverse conditions simultaneously. Continual learning methods [96, 97] incorporate strategies from the field of continual learning to effectively handle conditions that arise sequentially. Weight fusion methods [98, 99] assign weights to all conditions and blend these weights cohesively to ensure comprehensive control over all conditions. Attention-based integration methods [100, 99] modify the attention map to adaptively position and prioritize different conditions. Guidance composition methods [101, 102] integrate the independent denoising results of each condition to achieve a unified output.

### v3

@@ II-D Video LLM
Following the success of Image LLMs, researchers start exploring the training of Video LLMs [50]. Typically, videos are viewed as sequences of image frames (some Video LLMs incorporate other modalities like audio or speech), so Video LLMs have a higher computational complexity. The challenge of collecting high-quality video datasets further complicates the training process, making early fusion architectures computationally exhaustive. As a result, almost all the existing Video LLMs adopt the alignment architectures.
@@ II-D1 Alignment-Architecture Video LLM
The video LLM architecture is similar to that of Image LLMs with alignment architectures. By sampling a fixed number of frames or using a fixed frames-per-second (FPS) rate, videos are reduced to a limited set of images. The visual embeddings of each image are then extracted using a visual encoder. These features are sequentially concatenated in the order of the frames and connected to the LLM via an alignment module. In earlier works, VideoChat [51] utilizes a Q-former structure as the alignment module, while VideoLLaMA [52] introduces an audio encoder and an audio Q-former to handle audio signals. Video-ChatGPT [53] takes a different approach by average-pooling each frame’s patch embeddings along the spatial and temporal dimensions before using a linear layer as the alignment module. Training Video LLMs also follow an “alignment then instruction tuning” strategy. While additional GPT-annotated or human-annotated video datasets are collected, image datasets can also be leveraged by treating images as single-frame videos.
Recent successful efforts focus on improving performance by refining the alignment module and scaling up the model and dataset sizes. For instance, VideoLLaMA2 [54] improves the alignment module to model the connections across temporal and spatial dimensions. It also gathers datasets for tasks such as captioning, classification, and question answering. Qwen2.5-VL [55] and InternVL3 [56] leverage diverse training data, including images, videos, and interleaved image–text pairs, to build powerful vision-language models.
@@ II-D2 Challenges and Limitations in Video LLM
Remark. Despite the advancements in Video LLMs, nearly all existing models rely on sampling frames and encoding them individually through image encoders. This approach may be favored due to several reasons: image encoders are less computationally intensive compared to video encoders, they offer better alignment with textual data, and they facilitate unification with Image LLMs. However, this methodology comes with a significant limitation. Specifically, the process of sampling frames can lead to the complete loss of information that occurs between sampled frames. As a result, these models fail to capture the continuous motion and trajectories of objects, which are essential for understanding dynamic scenes and activities within a video.
II-E Speech LLM
Similar to Image LLMs, the architecture of Speech LLMs can generally be categorized into two types: alignment-based architectures and early-fusion architectures [64].
II-E1 Alignment-Architecture Speech LLM
This architecture first extracts information from audio with pre-trained or fine-tuned audio encoder and produces audio embedding.
a) Audio Encoder transforms raw waveforms into time–frequency representations using conventional signal processing techniques. The most commonly used audio encoders are Whisper [65] and Conformer [66]. Whisper is an automatic speech recognition (ASR) model with an encoder–decoder Transformer architecture, similar to sequence-to-sequence models in natural language processing. It is trained on 680,000 hours of multilingual, multitask supervised data collected from the web, covering speech recognition, speech translation, and language identification. Conformer (Convolution-augmented Transformer) combines convolutional neural networks (CNNs) with Transformer blocks, effectively capturing both local and global dependencies in speech signals. Other widely adopted encoders include WavLM [67], a self-supervised speech representation model built on the HuBERT [68] framework, with improvements in pretraining objectives and data diversity.
b) Alignment Module also referred to as a projector, connector, or adapter, maps audio embeddings into the text embedding space, enabling them to be processed by the LLM decoder for downstream understanding tasks. Several types of alignment modules have been proposed. One common approach is a multi-layer perceptron (MLP), which performs a straightforward projection. Another is the Q-Former, which introduces trainable query tokens that attend to audio features and produce fixed-length embeddings compatible with the LLM input space. A third approach is cross-attention, which allows bidirectional interactions between audio and text features, facilitating richer multimodal integration.
II-E2 Early-fusion Architecture Speech LLM
This type of Speech LLMs is inspired by visual tokenizers and adopts a similar approach for audio. In this framework, raw audio is converted into a sequence of discrete tokens that capture the acoustic content and can often be decoded back into high-quality audio. The generation of discrete tokens relies on vector quantization (VQ). Building on VQ-VAE [69], which introduced the idea of encoding continuous audio features into symbolic representations via a learned codebook, modern approaches include self-supervised pre-trained audio tokenizers such as HuBERT [68] and neural codec models such as EnCodec [70]. Several representative works fall under this branch of Speech LLMs. VALL-E [71] leverages EnCodec tokens to achieve zero-shot speech synthesis. SpeechGPT [72] is trained on paired unit-text data, where spoken audio is represented as discrete speech units. AudioPaLM [73] integrates wav2vec-style audio tokenization with language modeling to improve multimodal speech understanding.
Now we have discussed the multi-modal LLM for understanding. Next, we will discuss another important topic of multi-modal generative AI, i.e., multi-modal diffusion models for generation.
@@ III Multi-modal Diffusion for Generation
Figure 4: Comparison among GAN, VAE, diffusion, and flow matching models.
@@ III-A Preliminaries
We will first introduce some preliminaries, including traditional generative models, i.e., generative adversarial networks (GANs) and Variational AutoEncoders (VAEs). We then introduce diffusion probabilistic modeling and present a comparison among GAN, VAE, diffusion, and flow matching models, as illustrated in Fig. 4.
@@ III-A1 Generative Adversarial Networks
The generative adversarial network (GAN) [74] is one of the earliest neural architectures designed to generate visual content such as images [75] and videos [76]. The main idea of GANs involves two networks: a generator $\mathcal{G}$ and a discriminator $\mathcal{D}$. Specifically, $\mathcal{G}$ aims to generate visual content from a noise vector $z$, while $\mathcal{D}$ is trained to distinguish between real visual samples $x$ and generated ones $\mathcal{G}(z)$. These two networks are trained in an adversarial manner: the generator tries to produce outputs that can fool the discriminator, and the discriminator strives to accurately classify real versus fake samples. The training process forms a min-max game, where the generator learns to generate increasingly realistic samples to deceive a progressively stronger discriminator. The two networks are mutually reinforcing, so the training objective is as follows:
$\underset{\mathcal{G}}{min}\;\underset{\mathcal{D}}{max}\;\mathbb{E}_{x\sim p_{x}}\log\mathcal{D}(x)+\mathbb{E}_{z\sim p_{z}}\log(1-\mathcal{D}(\mathcal{G}(z))),$ (2)
where $z$ is sampled from $p_{z}$ that is usually a normal distribution and $x$ is a sample from the real data distribution $p_{x}$.
@@ III-A2 Variational AutoEncoder
Variational AutoEncoder [77] (VAE) is another typical generative model. Unlike GANs, autoencoders have an encoder-decoder architecture that uses an encoder $\mathcal{E}$ to present the visual content $x$ to a latent code $z=\mathcal{E}(x)$ and a decoder $\mathcal{D}$ to reconstruct the data $\hat{x}=\mathcal{D}(z)\approx x$. However, normal autoencoders have no constraints on the latent space, which makes them overfit the dataset easily. To solve the problem, VAEs make a regularization to the latent space and sample $z$ from a distribution $p_{\theta}$, typically a Gaussian distribution, where $\theta$ is the parameters of the encoder-decoder model. As the distribution $p_{\theta}$ is unknown, VAE utilizes a recognition model $\phi$ which serves as a variational approximation $q_{\phi}$ to approximate $p_{\theta}$ and trains them jointly:
$\mathcal{L}(\theta,\phi;x)=-D_{KL}(q_{\phi}(z|x)||p_{\theta}(z))+\mathbb{E}_{q_{\phi}(z|x)}[\log p_{\theta}(x|z)],$ (3)
where $D_{KL}$ means the Kullback-Leibler divergence. $\phi$ can be formulated as a differentiable estimator using the parameterization trick. To better generate visual content, many efforts [78, 69, 79] have been made based on VAE. Sync-DRAW [78] introduces a novel architecture that combines VAE with a recurrent attention mechanism to create a unique temporally dependent sequence of frames.
@@ III-A3 Diffusion Probabilistic Modeling
Remark. GANs, VAEs, and diffusion models are all generative models. Compared to GANs, which train both the generator and discriminator, the diffusion models have explicit probabilistic modeling and only train a denoising network $\epsilon_{\theta}$, which is more stable. Similarly, VAEs train both an encoder and a decoder. Moreover, diffusions denoise for each image $T$ times in the training phase, resulting in $T$ variants of each image as augmentation. These augmented images in turn help the denoising network to better model the data distribution $p_{\theta}(x_{0})$, leading to better generation results.
@@ III-A4 Latent Diffusion Model
Note that there is an additional input $c$ of the denoising network that is for conditional generation, e.g., as for the text-to-image generation task, $c$ could be the representation of the text prompt [82]. Also, $c$ could be other conditions, such as layout [83] or semantic maps [84]. Since most computation, including the training and iterative inference, is conducted in the lower-dimension latent space, the LDM model exhibits high efficiency. Therefore, most text-to-image and text-to-video models adopt the LDM structure.
III-A5 Flow Matching
Compared with diffusion models such as DDPM, Flow Matching [85] represents a new paradigm in generative modeling, built upon Continuous Normalizing Flows (CNFs). It introduces a simple yet intuitive training objective that learns to approximate a target vector field, which defines a probability path transforming noise samples into data samples. In this way, diffusion processes can be viewed as special cases within the broader Flow Matching framework.
Let $x_{1}$ denote a random variable drawn from an unknown data distribution $q(x_{1})$. We define a probability path $p_{t}$ such that $p_{0}=p$ is a simple distribution, e.g., the standard normal distribution $p(x)=\mathcal{N}(x|0,I)$, and $p_{1}$ approximates the data distribution $q$. The goal of Flow Matching is to learn a vector field that aligns the model’s probability path with this target path from $p_{0}$ to $p_{1}$.
$\mathcal{L}_{\rm FM}(\theta)=\mathbb{E}_{t,p_{t}(x)}\|v_{t}(x)-u_{t}(x)\|^{2},$ (6)
where $p_{t}(x)$ denotes the target probability density path, $u_{t}(x)$ is the corresponding vector field, and $v_{t}(x,\theta)$ is the learnable CNF vector field parameterized by $\theta$. Here $t\sim\mathcal{U}[0,1]$ is the uniform distribution, and $x\sim p_{t}(x)$. In essence, the Flow Matching loss trains the neural vector field $v_{t}$ to regress toward the target field $u_{t}$. When the loss approaches zero, the learned CNF model successfully reproduces the probability path $p_{t}(x)$.
Figure 5: Comparison between pixel-level and latent diffusion models.
@@ III-B Text-to-Image Generation
As mentioned in the preliminary part, diffusion models can be broadly categorized into two branches: pixel-based and latent-based [86]. In the early development stage, the denoising process is typically applied directly in the pixel space. For instance, GLIDE [87] is a pioneering work in photorealistic image generation with text guidance, using a 3.5 billion parameter diffusion model that employs a text encoder to condition on natural language descriptions. GLIDE also explores the use of CLIP guidance and classifier-free guidance in diffusion models, finding that classifier-free guidance produces higher-quality images. Besides, Imagen [88] follows GLIDE and adopts classifier-free guidance for its pixel-based diffusion model. The key difference between them is that GLIDE trains a text encoder and a diffusion model together with text-image pairs, while Imagen utilizes pretrained and frozen large transformer language models, leveraging their strong text understanding capabilities to enhance sample fidelity and image-text alignment.
@@
GLIDE [87], Imagen [88], Stable Diffusion [27], and DALL-E2 [89] are all pioneering works that represent different technological pathways in the field of text-to-image generation. These models have greatly inspired subsequent research and development [90, 91, 92]. Despite their differences, some common trends have emerged in their development. First, latent-based diffusion methods have become increasingly prevalent due to their advantages in conserving computational resources and generating high-quality images. Second, compared to classifier guidance [93], classifier-free guidance [94] is widely adopted in these works, where the label in a class-conditional diffusion model is replaced with a null label at a fixed probability during training. Third, U-Net traditionally serves as the backbone of the diffusion model, facilitating denoising and the gradual generation of high-quality images.
Despite its advantages in high-resolution image generation, U-Net’s specific structures, such as ResBlocks and convolutional operations, limit its scalability. In contrast, Transformers, which are better suited to handle larger-scale data and tasks, are emerging as strong contenders to U-Net. The Diffusion Transformer (DiT) [95] represents a class of diffusion models that replaces the commonly used U-Net backbone with a transformer backbone, as shown in Fig. 6. This approach is supported by empirical findings suggesting that the U-Net inductive bias is not crucial to the performance of diffusion models. Additionally, utilizing a transformer backbone enables the diffusion model to leverage the best practices of transformers, such as architectural design and training paradigms, along with their good properties, such as scalability, robustness, and efficiency. Specifically, DiT adheres to the foundation of the Latent Diffusion Model (LDM) framework and emulates the design of the Vision Transformer (ViT) by introducing a comprehensive DiT design space, including patch size, transformer block architecture, and model size. The first layer of DiT, termed patchify, converts the spatial input into a sequence of tokens by linearly embedding each patch. Following the patchify step, the input tokens are processed through a sequence of transformer blocks that incorporate conditioning, such as time and label. The proposed transformer design includes adaptive layer norm (adaLN) block, cross-attention block, and in-context conditioning block. After the final block, a transformer decoder translates the image tokens into output predictions. The difference between U-Net-based and Transformer-based diffusion models is illustrated in Fig. 6.